### v2

# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

Abstract
Open-set face forgery detection poses significant security threats and presents substantial challenges for existing detection models. These detectors primarily have two limitations: they cannot generalize across unknown forgery domains and inefficiently adapt to new data. To address these issues, we introduce an approach that is both general and parameter-efficient for face forgery detection. Our method builds on the assumption that different forgery source domains exhibit distinct style statistics. Specifically, we design a forgery-style mixture formulation that augments the diversity of forgery source domains, enhancing the model’s generalizability across unseen domains. In addition, previous methods typically require fully fine-tuning pretrained networks, consuming substantial time and computational resources. Drawing on recent advancements in vision transformers (ViT) for face forgery detection, we develop a parameter-efficient ViT-based detection model that includes lightweight forgery feature extraction modules and enables the model to extract global and local forgery clues simultaneously. We only optimize the inserted lightweight modules during training, maintaining the original ViT structure with its pre-trained ImageNet weights. This training strategy effectively preserves the informative pre-trained knowledge while flexibly adapting the model to the task of Deepfake detection. Extensive experimental results demonstrate that the designed model achieves state-of-the-art generalizability with significantly reduced trainable parameters, representing an important step toward open-set Deepfake detection in the wild.
Index Terms:
Deepfakes, face forgery detection, open-set, style mixture, generalization, robustness, parameter-efficient learning.
I Introduction
Figure 1: False Positive Rates (FPR) and False Negative Rates (FNR) of Xception, EfficientNet-B4, and ViT-B on four unforeseen Deepfake datasets: DFDC, DFR, WDF, and FFIW.
The rapid advancement of Artificial Intelligence Generated Content (AIGC) has led to significant progress in face forgery techniques, producing increasingly sophisticated fake faces that are indistinguishable from the human eye. Such advancements have fueled the proliferation of multimedia content on social media, posing high risks of disinformation, misinformation, and identity theft [1, 2]. The recent emergence of diffusion and other advanced large vision-language models has exacerbated these challenges [3]. Consequently, developing effective detection methods to combat digital face forgery attacks has become critically important [4].
Face forgery detection has long posed a challenging yet crucial problem. Early methodologies proposed extracting hand-crafted features such as LBP maps [5], color components [6], and DCT maps [7] as informative indicators for Deepfake detection. Subsequent works focused on capturing anomalies like eye-blinking [8], head pose artifacts [9], and face-warping features [10]. Many methods aimed to improve detection accuracy by developing learning-based models using generic architectures such as Xception Net [11], Efficient Net [12], and Capsule Net [13]. Later approaches, including Face X-ray [14], SBI [15], DCL [16], and RECCE [16] sought to capture common forgery features to enhance model generalizability. With the advent of ViT [17], numerous ViT-based detectors [18, 19, 20, 21] have been developed. These approaches exploit global receptive fields and exhibit superior detection performance. However, previous face forgery detection methods still face two major challenges in real-world applications: (1) Existing methods struggle to generalize in uncontrolled environments. These models may encounter unforeseen forgery domains in the wild, which can lead to significant performance degradation; (2) In real-world applications, detection models typically need fine-tuning with newly acquired data to adapt to different domains. However, this adaptation process for pre-trained detection models can be computationally expensive, making it challenging to deploy them effectively, particularly on mobile devices with limited computational resources.
Figure 2: Prediction score distributions of real and fake faces of four unseen Deepfake datasets: DFDC, DFR, WDF, and FFIW. The three rows represent three Deepfake detectors: Xception, Efficient Net, and ViT-B. It can be observed that domain gaps between different Deepfake datasets primarily impact the detection of forgery faces rather than real faces.
To address these two challenges, this paper introduces a generalized and parameter-efficient approach for Open-Set DeepFake Detection (OSDFD). Different forgery-style statistics may exhibit unique properties critical for face forgery detection. Fig. 1 shows that existing forgery detectors have a significantly higher false negative rate (FNR) than the false positive rate (FPR) in unforeseen forgery domains, indicating a tendency to misclassify fake faces as real when encountering new domains in uncontrolled environments.
In Fig. 2, we further demonstrate that domain gaps between different Deepfake datasets primarily impact the detection of forgery faces rather than real faces. Based on this observation, we propose a forgery style mixture module to increase the diversity of source forgery domains during training, thereby effectively mitigating forgery domain gaps in cross-dataset evaluations. Specifically, Fig. 2 presents cross-dataset experimental results for three representative face forgery detectors: Xception [11], EfficientNet [12], and ViT-B [17]. These models are trained on FF++ [22] and evaluated on DFDC [23], DFR [24], WDF [25], and FFIW [26]. In Fig. 2, green and red indicate real and fake faces, corresponding to ground-truth labels of ‘0’ and ‘1’, respectively. The $x$-axis and $y$-axis represent model prediction scores and the number of corresponding faces. As shown in Fig. 2, while these detectors perform well in identifying real faces, they struggle with fake faces in cross-dataset evaluations. This finding suggests that discrepancies between source and target datasets primarily arises from forgery domain gaps. Motivated by this observation, we design the forgery style mixture module to enhance the diversity of source forgery domains, thereby improving the model’s generalizability.
Furthermore, most prior art utilizes generic CNNs and ViTs as backbones and fully fine-tune the networks from ImageNet weights. However, fully fine-tuning a network is costly, hindering its real-world applications. To mitigate this issue, we integrate lightweight Adapter and LoRA layers into the ViT backbone. During training, we freeze plain ViT backbones with ImageNet weights and solely update the lightweight Adapter and LoRA layers. These designed modules capture local and global forgery artifacts using slim components. Additionally, these modules preserve substantial ImageNet knowledge and enable flexible model adaptation for face forgery detection. While parameter-efficient fine-tuning (PEFT) has been widely adopted in various domains, our work makes a notable contribution by integrating PEFT with a forgery-aware perspective, specifically tailored for open-set face forgery detection—a task that presents unique challenges in generalizability and adaptation. Fig. 3 illustrates the average cross-dataset AUC performance across six unforeseen Deepfake datasets versus the number of activated parameters. This demonstrates that our method achieves superior generalizability with the fewest activated parameters.
This work extends upon our previous conference work [27]. The main distinctions between this work and [27] are summarized as follows: (1) We introduce the Central Difference Convolution (CDC) adapter to incorporate local forgery priors into the expressive ViT model, thereby enhancing the model’s detection performance, (2) We propose a forgery style mixture module that combines source forgery domains to improve the model’s generalizability across unforeseen forgery domains; (3) We further reduce the dimension of LoRA layers, enabling more parameter-efficient face forgery detection; (4) We conduct extensive additional experiments compared with more SOTA methods, including cross-dataset evaluations, robustness experiments, visualization results, and ablation studies.
Figure 3: Average cross-dataset AUC score across six unseen datasets vs. the number of activated parameters. Our OSDFD method achieves the best generalizability with the fewest trainable parameters.
In summary, the main contributions of this work are as follows:
  • We design a forgery-style mixture module to enhance the diversity of source domains. The model’s generalizability is effectively improved by randomly mixing forgery styles.
  • Our approach explicitly captures forgery-related artifacts by integrating a CDC adapter and a LoRA module into a ViT-based detection framework. This targeted design enhances the model’s ability to extract both global and local forgery cues, which is crucial for Deepfake detection.
  • Our designed components jointly contribute to an effective and scalable solution for open-set Deepfake detection. Extensive experiments demonstrate that our model achieves state-of-the-art generalizability and robustness with only 1.34M activated parameters.
We arranged the remaining of this paper into four more sections, Sec. II comprehensively reviews previous literature on face forgery detection and parameter-efficient fine-tuning. Sec. III details the proposed detection framework. Sec. IV presents extensive experimental results under various settings—finally, Sec. V concludes the paper and discusses possible future research directions.
II Related Work
II-A Face forgery detection
Early face forgery detection methods initially focused on extracting hand-crafted features, such as eye-blinking frequency [8], head pose inconsistency [9], face warping artifacts [10], and rPPG features [28]. With the rise of deep learning and artificial intelligence, numerous learning-based methods emerged to improve detection accuracy. Some generic CNNs, such as the Xception network [11, 29, 30, 7, 31], EfficientNet-b4 [12, 32, 33, 34], and Capsule network [13], have been utilized for face forgery detection, showing promising results. However, these methods often need to be more balanced in training data, leading to poor generalization capabilities. In response, some approaches have focused on extracting common forgery features. For example, Face X-ray [14] and SBI [15] aimed to extract common face blending features, enhancing model generalizability. F3Net and Luo et al. [30] mined frequency artifacts, while DCL [16] and CFM [32] employed contrastive learning schemes for more generalized Deepfake detection. NoiseDF [35] raised an innovative idea to distinguish Deepface-specific noise traces from common noises within each single frame, providing additional interpretability. With the rapid advancements in vision transformers (ViT), several ViT-based approaches [21, 18, 27, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45] have been developed, showing superior detection performance. ICT [21] introduced the Identity Consistency Transformer to capture identity inconsistency within fake faces, while F2Trans designed the High-Frequency Fine-Grained Transformer, achieving more generalized and robust forgery detection. However, the scalability of ViT-based methods is largely constrained by the quadratic complexity of attention computation, limiting their deployment, particularly on mobile devices with limited computing resources. In response, this paper introduces a parameter-efficient fine-tuning (PEFT) strategy for face forgery detection to address this issue. Furthermore, we combine different forgery styles to augment the forgery source domains, enhancing the model’s generalization capability.
II-B Parameter-efficient fine-tuning (PEFT)
The attention mechanism has demonstrated significant success across various computer vision tasks. However, the self-attention in ViT scales quadratically concerning the input patch number, leading to substantial computational demands during training. Several methods have been proposed to address this issue, including parameter-efficient fine-tuning (PEFT). PEFT selectively updates a small portion of the model parameters, facilitating model adaptation to downstream tasks. Adapters [46, 47] are typically inserted between feedforward and normalization layers of transformers, comprising a down-sample and up-sample layer. For example, Low-Rank Adaptation (LoRA) [48] decomposes a matrix into two trainable low-rank matrices, significantly reducing the number of parameters compared to the original matrix. Visual Prompt Tuning (VPT) [49] introduces a small number of trainable parameters in the input space, treating additional tokens as learnable pixels. Convpass [50] aims to inject local priors into expressive ViT models by integrating convolutional operations, enhancing performance across various computer vision tasks. Neural Prompt Search (NOAH) [51] further advances by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Despite these advancements, the development of PEFT methods for face forgery detection has been largely understudied, yet it is crucial in practical scenarios. In this paper, we integrate lightweight Adapter and LoRA modules into the plain ViT backbone, optimizing only these modules while keeping the ViT backbone fixed with ImageNet weights during training. The designed modules effectively capture global and local face forgery features more efficiently, without the danger of catastrophic forgetting typical of fully finetuning models [52].
Figure 4: Overview of the designed face forgery detection framework. (a) Overall model structure; (b) Structure of the designed transformer block; (c) Details of the designed adapter layer; (d) Details of the designed LoRA layer.
Figure 5: Illustration of the Central-Difference Convolution (CDC) pipeline.
III Proposed Method
III-A Overall framework
Fig. 4 illustrates the overview of the proposed face forgery detection method. The designed framework relies upon ViT, which is pre-trained on ImageNet as the backbone. As Fig. 4 (a) depicts, the features extracted by the transformer blocks are fed forward to the Forgery Style Mixture module. This module exclusively mixes the statistics of forgery styles, augmenting the forgery source domains during training. The processed features are delivered to the MLP head for the final decision-making. Fig. 4 (b) illustrates the details of transformer blocks. We insert the trainable Adapter and LoRA layers into the ViT block. The ViT backbone is fixed with pre-trained ImageNet weights. Fig. 4 (c) and Fig. 4 (d) detail the Adapter and LoRA layers. In Fig. 4 (c), the input hidden state $h^{in}$ is followed by three trainable convolutional layers. Moreover, $h^{in}$ is simultaneously multiplied with the ImageNet pretrained weight. The output hidden state $h^{out}$ is the element summation of the two features. In turn, the LoRA layer dedicated in Fig. 4 (d) introduces two trainable low-rank matrices $W^{down}\in\mathbb{R}^{d\times r}$ and $W^{up}\in\mathbb{R}^{r\times d}$, where $r<<d$. Moreover, $h^{out}$ is obtained by conducting an element-wise summation of the two outputs.
Figure 6: Illustration of forgery style mixture pipeline. The designed feature extractor processes real faces and forgery source domain data, and the extracted features are subsequently delivered to the forgery-style mixture module.
III-B Forgery-aware PEFT
Our proposed method aims to capture critical forgery clues using lightweight modules to facilitate face forgery detection in practical scenarios. We integrate forgery-aware PEFT modules into plain ViTs. These modules include the LoRA layer, injected into the self-attention blocks, and the adapter layer, integrated into the feed-forward networks (FFN).
As Fig. 4 (c) depicts, the input hidden state $h^{in}$ is simultaneously forwarded to the fixed ImageNet MLP weights and the designed trainable convolutional layers. Subsequently, the outputs of these processes are element-wise summed before being passed to the subsequent vision transformer blocks:
$h^{out}={\rm MLP}(h^{in})+{\rm Adapter}(h^{in}).$ (1)
The Adapter layer is formulated as:
${\rm Adapter}(h^{in})={\rm Conv_{1\times 1}^{up}}({\rm CDC}({\rm Conv_{1\times 1}^{down}}(h^{in}))).$ (2)
We first expand the input hidden state $h^{in}$ into a 2-D dimension. Next, we reduce the dimension of the expanded feature using a $1\times 1$ convolutional layer, effectively decreasing the number of parameters activated in the Adapter layer. We then design the Central Difference Convolution (CDC) operator to expose local forgery artifacts for each input query. The pipeline of the CDC operator is illustrated in Fig. 5. The output feature map $x^{out}$ can be calculated by:
$x^{out}=\sum_{i\in\Omega}w_{i}(x^{in}_{i}-x_{c}),$ (3)
where $x_{c}$ indicates the center pixel of the window $\Omega$, and $x_{5}$ represents $x_{c}$ in Fig. 5. $x_{i}$ indicates the peripheral pixels within $\Omega$. $w_{i}$ denotes the corresponding convolutional weights. Finally, the second $1\times 1$ Conv layer restores the channels of the feature, and the produced feature is flattened to the same shape as $h^{in}$. Our designed adapter layer effectively introduces inductive bias into standard ViTs, guiding the model in learning local artifacts for face forgery detection. Additionally, considering the effectiveness of high-frequency features in the task of face forgery detection [7, 53, 54, 55], our CDC operator enables the model to extract abundant high-frequency information from input faces, thereby enhancing detection performance.
In addition to the Adapter layer, which introduces local dependencies into plain ViT backbones, we further design the Low-Rank Adaptation (LoRA) layer to obtain global receptive fields with minimal parameters. As depicted in Fig. 4 (d), the output hidden state $h^{out}$ is calculated by:
$h^{out}=W_{q/k/v}h^{in}+{\rm LoRA}(h^{in}).$ (4)
Figure 7: Pipeline of the proposed forgery feature mixture process.
The $W_{q/k/v}$ parameter denotes the pre-trained ImageNet weights fixed during the training process. Moreover, the LoRA module consists of two low-rank matrices:
${\rm LoRA}(h^{in})=h^{in}W_{q/k/v}^{down}W_{q/k/v}^{up},$ (5)
where $h^{in}\in\mathbb{R}^{d\times d}$, $W_{q/k/v}^{down}\in\mathbb{R}^{d\times r}$, $W_{q/k/v}^{up}\in\mathbb{R}^{r\times d}$, and $r<<d$. We optimize the slim, decomposed low-rank matrices during training instead of the redundant pre-trained weights. As such, the number of trainable parameters is significantly reduced. In this work, we set the hidden state dimension $d$ of ViT-B to 768, and the rank $r$ is defaulted to 8. The designed LoRA and CDC adapter respectively capture expressive global features and extract informative local forgery clues, thereby enhancing the final detection performance.
The designed forgery-aware PEFT modules bear the following advantages for face forgery detection: (1) By optimizing a small subset of parameters, our method reduces computational and memory requirements compared to full fine-tuning while also accelerating training, making frequent updates more practical in open-set environments. Moreover, storing the backbone weights in the cloud while keeping only a minimal set of parameters on edge devices facilitates deployment on resource-constrained mobile platforms. (2) As model sizes continue to grow, our forgery-aware PEFT strategy remains scalable for large models, and the designed modules can be seamlessly integrated with various backbones in a plug-and-play fashion. (3) Our approach efficiently incorporates forgery-specific knowledge while preserving the pretrained model’s general knowledge, effectively mitigating overfitting to forgery data and enhancing generalization performance.
III-C Forgery style mixture
In face forgery detection, the training dataset often encompasses multiple forgery techniques. This training setting accommodates practical scenarios where available source domains are utilized. In this vein, we denote the training dataset as $D^{S}=\{R^{S},F_{1}^{S},F_{2}^{S},...,F_{N}^{S}\}$. $R^{S}$ represents the real source faces, while $F_{i}^{S}$ denotes the $i_{th}$ source forgery domain, where $i\in\{1,2,...,N\}$. Following previous methods in face forgery detection, we use FF++ as the training set, which includes multiple training forgery domains (i.e., manipulation types). As such, the target testing domain can be denoted as $D^{T}=\{R^{T},F^{T}\}$. Similarly, $R^{T}$ and $F^{T}$ are target real and forgery domains. Our preliminary experiments, illustrated in Fig. 1 and Fig. 2, reveals a high false negative rate (FNR) and a low false positive rate (FPR) in cross-domain evaluations, suggesting a negligible domain gap between $R^{S}$ and $R^{T}$ and significant domain gaps between $F^{S}$ and $F^{T}$. As such, we propose to augment source forgery domains to mitigate this issue. Fig. 6 depicts the proposed forgery-style mixture pipeline. Each training batch, composed of one real domain and multiple fake domains, is processed by our proposed feature extractor, which shares weights across all source domains. The extracted features from the source forgery domains are passed to the designed forgery-style mixture module, and the mixed features are subsequently forwarded to the classifier to perform forgery detection.
Next, we elaborate on the details of the devised forgery style mixture module, as depicted in Fig. 7. Green squares denote real faces in this representation, while red circles indicate the forgery faces. This module processes each training batch consisting of real and fake face images sampled from the real source domain and $N$ forgery source domains. In Fig. 7, the subscripts of $f$ represent the domain labels, while the superscripts denote the sample index within each domain. We illustrate the case of NN=3 forgery domains in Fig. 7, with 3 samples per domain. The features $\mathcal{F}_{in}$ of each input batch is proceeded to our designed Forgery Style Mixture module.
We first sort $\mathcal{F}_{in}$ to $\mathcal{F}_{sort}$ according to the domain labels, where $f_{R}$ represents real features, and $f_{Fi}$ with $i\in\{1,2,3\}$, denotes features from different forgery domains. To augment the forgery style, we first shuffle the fake features of $\mathcal{F}_{sort}$, ensuring that each forgery domain label of the shuffled $\widetilde{\mathcal{F}}_{sort}$ differs from the original $\mathcal{F}_{sort}$. In addition, we keep the real feature order unchanged. Subsequently, we mix the original $\mathcal{F}_{sort}$ and the shuffled $\widetilde{\mathcal{F}}_{sort}$ by fusing their feature statistics. The details of this process are provided next.
We draw inspiration from Adaptive Instance Normalization (AdaIN) [56], a widely employed image style transfer technique. The source style of $p$ can be transferred to the target style of $q$ by:
${\rm AdaIN}(p)=\sigma(q)\frac{p-\mu(p)}{\sigma(p)}+\mu(q),$ (6)
where $\mu(p)$ and $\sigma(p)$ are the spatial mean and standard deviation of the feature tensor $p$.
Inspired by the AdaIN, which utilizes the feature statistics of a target style input to transfer the image style, we first calculate the mixed feature statistics of $\mathcal{F}_{sort}$ and $\widetilde{\mathcal{F}}_{sort}$:
[…]
where $\delta$ is a weight sampled from Beta distribution, i.e., $\delta\sim Beta(0.1,0.1)$. $\gamma_{mix}$ and $\eta_{mix}$ are the weighted summation of the forgery style statistics of $\mathcal{F}_{sort}$ and $\widetilde{\mathcal{F}}_{sort}$. Then, the style mixture is performed by the mixed feature statistics $\gamma_{mix}$ and $\eta_{mix}$ to the original feature $\mathcal{F}_{sort}$:
[…]
Finally, we restore the original index order of the mixed features $\mathcal{F}_{mix^{\prime}}$ to obtain $\mathcal{F}_{mix}$, which is forwarded to the classifier for final decision-making. Similar to [57], we set a default probability of 0.5 to activate the forgery-style mixture module. In the inference stage, the forgery-style mixture is deactivated. As such, the designed module augments the source forgery domains in the training phase, thereby mitigating the overfitting problem and enhance the model’s generalizability. This design effectively improves the model’s ability to generalize in open-set face forgery detection.
TABLE I: High quality (c23) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
[…]
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet18 [58] CVPR16 11.18M 0.813 0.656 0.746 0.596 0.464 0.476 0.688 0.528 0.678 0.564
Xception [11] CVPR17 20.81M 0.907 0.795 0.753 0.558 0.460 0.472 0.744 0.557 0.716 0.596
EN-B4 [12] ICML19 19.34M 0.485 0.495 0.556 0.523 0.517 0.517 0.493 0.500 0.513 0.509
All-train EN-B4 [12] ICML19 19.34M 0.911 0.824 0.801 0.633 0.543 0.500 0.774 0.608 0.757 0.641
Focal-loss EN-B4 [12] ICML19 19.34M 0.903 0.813 0.798 0.608 0.503 0.484 0.759 0.604 0.741 0.627
Forensics Transfer [59] ArXiv18 - - 0.720 - 0.645 - 0.460 - 0.569 - 0.599
Multi-task [60] BTAS19 - - 0.703 - 0.587 - 0.497 - 0.603 - 0.598
MLDG [61] AAAI18 62.38M 0.918 0.842 0.771 0.634 0.609 0.527 0.780 0.621 0.770 0.656
LTW [62] AAAI21 20.37M 0.927 0.856 0.802 0.656 0.640 0.549 0.773 0.653 0.786 0.679
ViT-B [17] ICLR21 85.80M 0.771 0.701 0.656 0.582 0.510 0.498 0.554 0.517 0.623 0.575
CFM [63] TIFS24 25.37M 0.880 0.617 0.814 0.557 0.631 0.516 0.644 0.508 0.742 0.550
DCL [16] AAAI22 19.35M 0.949 0.877 0.829 0.684 - - - - - -
Implicit ID [64] CVPR23 22.36M 0.950 0.882 0.844 0.694 - - - - - -
UCF [65] CVPR23 20.81M - 0.867 - 0.679 - - - - - -
DiffusionFake [66] NIPS24 19.34M - 0.882 - 0.702 - - - - - -
SCLoRA [27] MIPR23 4.72M 0.935 0.862 0.875 0.753 0.651 0.554 0.707 0.626 0.792 0.699
OSDFD Ours 1.34M 0.962 0.885 0.932 0.846 0.698 0.595 0.781 0.682 0.843 0.752
TABLE II: Low quality (c40) cross-manipulation detection performance on unseen forgery methods. The trainable parameter number of each model is reported to evaluate the training efficiency.
[…]
AUC ACC AUC ACC AUC ACC AUC ACC AUC ACC
ResNet18 [58] CVPR16 11.18M 0.741 0.673 0.648 0.600 0.634 0.594 0.598 0.567 0.655 0.609
Xception [11] CVPR17 20.81M 0.766 0.694 0.696 0.643 0.626 0.593 0.597 0.552 0.671 0.621
EN-B4 [12] ICML19 19.34M 0.451 0.485 0.537 0.505 0.512 0.503 0.499 0.497 0.500 0.498
All-train EN-B4 [12] ICML19 19.34M 0.753 0.676 0.674 0.614 0.614 0.580 0.600 0.564 0.660 0.609
Focal-loss EN-B4 [12] ICML19 19.34M 0.749 0.674 0.672 0.610 0.596 0.575 0.605 0.566 0.656 0.606
Forensics Transfer [59] ArXiv18 - - 0.682 - 0.550 - 0.530 - 0.550 - 0.578
Multi-task [60] BTAS19 - - 0.667 - 0.565 - 0.517 - 0.560 - 0.577
MLDG [61] AAAI18 62.38M 0.730 0.671 0.617 0.581 0.617 0.581 0.607 0.569 0.643 0.601
LTW [62] AAAI21 20.37M 0.756 0.691 0.724 0.657 0.681 0.625 0.608 0.585 0.692 0.640
ViT-B [17] ICLR21 85.80M 0.739 0.643 0.650 0.595 0.592 0.560 0.552 0.538 0.633 0.584
CFM [63] TIFS24 25.37M 0.789 0.686 0.622 0.541 0.688 0.607 0.566 0.517 0.666 0.588
DCL [16] AAAI22 19.35M 0.838 0.759 0.751 0.679 - - - - - -
Implicit ID [64] CVPR23 22.40M 0.846 0.769 0.748 0.680 - - - - - -
UCF [65] CVPR23 20.81M - 0.746 - 0.673 - - - - - -
DiffusionFake [66] NIPS24 19.34M - 0.751 - 0.713 - - - - - -
SCLoRA [27] MIPR23 4.72M 0.818 0.735 0.686 0.638 0.710 0.653 0.582 0.542 0.699 0.642
OSDFD Ours 1.34M 0.853 0.762 0.717 0.663 0.713 0.661 0.630 0.595 0.728 0.670
III-D Objective function
During training, as we augment source-forgery features using the designed forgery style mixture module, we must train a classifier that can effectively distinguish between real face features and the augmented forgery features. Therefore, this work adopts the Single-Center Loss (SCL) [62] to obtain a clearer boundary between bonafide and fake faces. The overall objective function consists of two losses:
$L=L_{BCE}+\lambda L_{SCL},$ (10)
where $\lambda$ balances the two loss components. $L_{BCE}$ represents the binary cross-entropy loss. The Single-Center Loss, denoted as $L_{SCL}$, aims to compact the feature distribution of real faces while simultaneously moving fake features away from the center of real features:
$L_{SCL}=Dist_{R}+max(Dist_{R}-Dist_{F}+margin,0)$ (11)
where $Dist_{R}$ and $Dist_{F}$ denote the average distance between the real feature center $C$ and real feature $f_{R}$ and fake feature $f_{F}$, respectively:
[…]
where the real center $C$ is the average of all $N_{R}$ real features $f_{R}$:
$C=\frac{1}{N_{R}}\sum_{{i=1}}^{N_{R}}f_{R}^{i}.$ (13)
We select the features after the second-to-last fully-connected layer to calculate $L_{SCL}$. This loss enhances the discriminative and separable characteristics of the real and fake face features, thereby improving the overall performance of face forgery detection.
IV Experiments and Results
IV-A Experimental setup
IV-A1 Implementation details
In this study, we employ the popular dlib face detector [67]. The detected face regions are enlarged by a factor of 1.3 and subsequently resized to 224×224 pixels before being input into the model. The proposed framework is implemented on the PyTorch [68] platform, and the model is trained using Adam optimizer [69] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate at 3e-5, and no learning rate decay is applied throughout the whole training process. Our default architecture employs the ViT-B as the backbone, and the ViT-B parameters are initialized with ImageNet-21k [70] weights. We set the batch size to 48, as a single 2080Ti GPU can accommodate a maximum of 48 images at a resolution of $\times224$. During training, we use AUC as the validation metric. The training process consists of 30,000 iterations, as we observe that the validation AUC converges after about 20,000 iterations.
IV-A2 Datasets
We adopt the FaceForensics++ (FF++) dataset [22] for training, which includes original real videos (OR) and four types of forgeries: Deepfakes (DF) [71], Face2Face (FF) [72], FaceSwap (FS) [73], and NeuralTextures (NT) [74]. Each category comprises 1,000 videos in raw, high-quality (c23), and low-quality (c40) formats. Following the official data split strategy [22], we allocate 720 videos for training, 140 for validation, and 140 for testing. Raw videos represent non-compressed videos, whereas c23 and c40 videos denote slightly compressed and highly compressed videos. Our experiments exclusively employ the compressed c23 and c40 videos to align with practical applications.
IV-A3 Evaluation Metrics
To facilitate head-to-head comparisons with previous methodologies, we employ Accuracy (ACC), Area Under the receiver operating characteristic Curve (AUC), and Equal Error Rate (EER) as the evaluation metrics in this work. ACC is a straightforward metric that quantifies the proportion of correctly classified samples from the total number of samples. AUC measures the Area under the Receiver Operating Characteristic (ROC) curve. EER denotes the point of ROC where False Positive Rate (FPR) is equal to the True Positive Rate (TPR). Unless specified otherwise, we report the image-level detection results throughout this paper.
IV-B Cross-manipulation evaluation
In real-world applications, face forgery detectors inevitably encounter data generated by new forgery techniques, leading to significant performance drops. To evaluate the generalization capability, we train the models on the training sets comprising three source manipulation types of FF++. Subsequently, we evaluate the trained models on the test sets of the unseen target manipulation type. Table I presents the detection AUC and ACC of different models on the FF++ high-quality (c23) dataset. We bold the best detection results and underline the second-best results. It can be observed that OSDFD achieves SOTA AUC and ACC performances across all experiments.
Moreover, generalization to poor-quality data is crucial for real-world face forgery detectors. Therefore, we conduct the cross-manipulation experiments on low-quality (c40) data, as shown in Table II, where OSDFD outperforms previous approaches by a significant margin. This underscores the proposed method’s outstanding generalizability across diverse data qualities. Additionally, our approach demonstrates substantial improvements over the ViT-B baseline, demonstrating the effectiveness of the designed framework. Furthermore, compared to our conference version SCLoRA [27], our method consistently achieves superior performance on both c23 and c40 data, which can be attributed to the effectiveness of the designed forgery style mixture module and local CDC feature extractor. Finally, OSDFD comprises the fewest 1.34M trainable parameters, facilitating its deployment in real-world applications, particularly on mobile devices.
TABLE III: Cross-dataset evaluation on six unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Methods highlighted in blue denote video-level results. The trainable parameter number of each model is reported to evaluate the training efficiency.
Method Venue #Params CDF WDF DFDC-P DFDC DFR FFIW Avgerage
AUC EER AUC EER AUC EER AUC EER AUC EER AUC EER AUC EER
Face X-ray [14] CVPR20 41.97M 74.20 - - - 70.00 - - - - - - - - -
GFF [30] CVPR21 53.25M 75.31 32.48 66.51 41.52 71.58 34.77 - - - - - - - -
LTW [62] AAAI21 20.37M 77.14 29.34 67.12 39.22 74.58 33.81 - - - - - - - -
F2Trans-S [75] TIFS23 117.52M 80.72 - - - 71.71 - - - - - - - - -
SBI* [15] CVPR22 19.34M 81.33 26.94 67.22 38.85 79.87 28.26 68.45 37.23 84.90 23.13 81.47 27.33 77.21 30.29
DCL* [16] AAAI22 19.35M 81.05 26.76 72.95 35.73 71.49 35.90 72.69 34.91 92.26 14.81 82.08 27.75 78.75 29.31
Xception†[11] ICCV19 20.81M 64.14 39.77 68.90 38.67 69.56 36.94 64.52 39.29 91.93 15.52 77.58 31.21 72.77 33.57
RECCE†[76] CVPR22 25.83M 61.42 41.71 74.38 32.64 64.08 40.04 67.61 36.98 92.93 14.74 85.84 24.12 74.38 31.71
EN-B4†[12] ICML19 19.34M 65.24 39.41 67.89 37.21 67.96 37.60 63.48 40.14 92.18 15.51 70.03 35.40 71.13 35.40
Implicit ID [64] CVPR23 22.36M 83.80 24.85 - - - - 81.23 26.80 - - - - - -
SFDG [77] CVPR23 22.40M 75.83 30.30 69.27 37.70 - - 73.64 33.70 92.10 15.10 - - - -
UCF* [65] CVPR23 46.80M 77.15 30.20 77.40 29.80 69.30 35.80 68.40 37.80 80.80 27.00 68.30 36.70 73.56 32.88
LSDA [78] CVPR24 20.81M 82.70 - 71.40 - 76.30 - 71.10 - 87.50 - 84.20 - 78.87 -
DiffusionFake [66] NIPS24 19.34M 83.17 24.59 75.17 33.25 77.35 30.17 - - - - - - - -
DSRL* [79] IJCV24 128.31M 74.23 30.48 74.18 31.16 72.01 34.03 73.74 31.35 94.50 13.09 86.77 20.41 79.24 26.75
FreqNet† [80] AAAI24 1.85M 72.81 33.01 73.37 32.67 72.89 33.37 68.78 36.65 86.94 26.48 75.67 30.26 75.08 32.07
CFM [63] TIFS24 25.37M 82.78 24.74 78.39 30.79 75.82 31.67 68.28 37.14 95.18 11.87 78.71 29.28 79.86 27.58
ViT-B [17] ICLR21 85.80M 72.35 34.50 75.29 33.40 75.58 32.11 71.48 35.54 80.47 26.97 68.49 37.12 73.94 33.27
SCLoRA [27] MIPR23 4.72M 73.56 32.87 77.91 29.56 77.08 30.37 72.32 34.52 94.78 11.76 81.42 26.59 79.51 26.59
OSDFD Ours 1.34M 83.35 24.47 78.53 28.46 76.28 31.17 75.52 31.99 95.54 9.90 84.61 24.09 82.31 25.01
Lisiam [81] TIFS22 - 78.21 - - - - - - - - - - - - -
F3Net [7] ECCV20 42.53M 68.69 - - - 67.45 - - - - - - - - -
FTCN [82] ICCV21 26.60M 86.90 - - - 74.00 - - - - - - - - -
SBI* [15] CVPR22 19.34M 88.61 19.41 70.27 37.63 84.80 25.00 71.70 35.27 90.04 17.53 86.34 23.45 81.96 26.38
DCL* [16] AAAI22 19.35M 88.24 19.12 76.87 31.44 77.57 29.55 75.03 30.94 97.41 9.96 86.26 24.32 83.56 24.22
RECCE†[76] CVPR22 25.83M 69.25 34.38 76.99 30.49 66.90 39.39 70.26 34.25 97.15 9.29 92.05 17.74 78.77 27.59
UCF* [65] CVPR23 46.80M 83.70 28.20 77.20 26.10 71.00 33.50 70.90 35.80 86.40 20.40 69.70 34.20 76.48 29.70
LSDA [78] CVPR24 20.81M 89.80 - 75.60 - 81.20 - 73.50 - 89.20 - 84.20 - 82.25 -
DSRL* [79] IJCV24 128.31M 84.56 22.90 77.31 27.32 75.64 30.17 77.99 27.63 97.23 10.44 88.69 19.82 83.57 23.05
FreqNet†[80] AAAI24 1.85M 75.79 28.59 79.54 29.14 75.07 31.16 69.51 36.11 88.79 25.97 79.56 26.47 78.04 29.57
CFM [63] TIFS24 25.37M 89.65 17.65 82.27 26.80 80.22 27.48 70.59 35.02 97.59 9.04 83.81 24.49 84.02 23.41
ViT-B [17] ICLR21 85.80M 78.12 30.59 78.95 29.59 79.43 28.62 73.35 33.05 86.42 19.90 70.34 35.45 77.77 29.53
SCLoRA [27] MIPR23 4.72M 81.11 26.32 82.78 24.85 80.20 26.68 74.96 35.01 98.53 6.63 85.59 22.70 83.86 23.70
OSDFD Ours 1.34M 90.81 16.94 81.42 24.90 79.49 28.31 78.74 29.37 98.89 4.48 90.01 17.50 86.56 20.25
IV-C Cross-dataset evaluation
Due to the significant domain gaps between training and testing datasets [83] in face forgery detection, cross-dataset evaluation has been a persistent challenge. We further conduct cross-destatet evaluations to assess the generalizability of forgery detectors. To ensure fair comparisons in cross-dataset experiments, we follow previous works [63, 76, 16] by using the FF++ (c23) dataset [22] as the training set. This dataset includes original real faces and four types of forgeries: Deepfakes, Face2Face, FaceSwap, and NeuralTextures. To balance positive and negative samples during training, real faces are augmented fourfold. Additionally, during the forgery style mixture stage, we randomly mix the four forgery types. We test the trained models on six popular Deepfake datasets, including CelebDF-v2 (CDF) [84], WildDeepfake (WDF) [25], DeepFake Detection Challenge (DFDC) [23], DeepFake Detection Challenge Preview (DFDC-P) [23], DeepForensics-1.0 (DFR) [24], and Face Forensics in the Wild (FFIW) [26]. These testing datasets simulate real-world scenarios well, encompassing various environmental variables. AUC and EER performance are reported in Table III, where ‘*’ indicates the trained model provided by the authors, and ‘†’ indicates our re-implementation using the public official code. The upper table presents the frame-level detection results, while the blue-highlighted bottom table shows video-level results. Video-level detection scores are calculated by averaging all frame scores within each video. The best and second-best results are highlighted in bold and underlined, respectively. The trainable parameter number of each detector is provided in the table for reference. Our method, OSDFD, achieves the highest average detection performance in both frame-level and video-level evaluation, with only 1.34M activated parameters. Compared to the ViT-B backbone, OSDFD consistently demonstrates significant performance improvements across all six datasets, demonstrating its exceptional generalization capability. Furthermore, OSDFD outperforms our conference version SCLoRA by a significant margin, underscoring the effectiveness of the designed CDC Adapter layer and forgery-style mixture module.
IV-D Robustness evaluation
We incorporate five severity levels for each perturbation type to mimic unpredictable environments. The robustness evaluation results are depicted in Fig. 8. OSDFD does not employ data augmentations during training, facilitating efficient training, or fine-tuning for real-world applications. All models depicted in Fig. 8 are trained on the FF++ c23 training set, with distortions applied to the corresponding test set. As the severity of distortion increases, all face forgery detection methods experience performance degradation. Compared to previous arts, our proposed approach achieves superior or comparable AUC performance across all six types of perturbation. Furthermore, the average robustness results across all distortions highlight that OSDFD, with the fewest trainable parameters, demonstrates the best robustness against practical image/video perturbations. Our proposed style mixture module contributes to enhancing model robustness by increasing the diversity of forgery domains, thereby reducing the distribution gap between clean training data and corrupted testing data.
Figure 8: Robustness to common perturbations at five severity levels: Gaussian Blur, Pink Noise, White Noise, Block, Brightness, and Darkness.
TABLE IV: Ablation Experiments on different model components. Average Cross-Manipulation detection performance on FF++ c23 dataset.
ViT LoRA Adapter MixStyle SCL AUC ACC
- - - - 0.623 0.575
- - - 0.816 0.719
- - - 0.825 0.723
- - - 0.795 0.712
- - 0.832 0.740
- 0.836 0.743
0.843 0.752
IV-E Ablation study
In this subsection, we carry out extensive ablation experiments to showcase the effectiveness of the designed components, the flexibility of the proposed learning scheme, the rationale behind the adopted training strategy, and the impacts of the mixed features.
IV-E1 Effectiveness of the designed modules
We conduct ablation experiments to study the impact of each designed module. Table IV reports the average cross-manipulation (c23) detection AUC and ACC scores across four experimental trials. The designed LoRA and Adapter modules significantly enhance the model’s generalizability. The combination of LoRA and Adapter enables the model to simultaneously explore global and local anomalies within input forgery face images, enhancing the final detection performance. The forgery-style mixture module further boosts the AUC and ACC scores by augmenting the forgery source domains during training. The Single-Center Loss (SCL) encourages the real features to cluster more tightly while pushing features from fake images away from the real center. This design leads to a clearer decision boundary between real and fake faces, further enhancing the generalizability of the detection. Overall, the experimental results presented in Table IV underscore the effectiveness of the designed components.
The proposed method showcases its adaptability by being flexibly adapted to other transformer backbones. OSDFD comprises forgery-aware, parameter-efficient fine-tuning modules and a forgery-style mixture module, both designed as plug-and-play components that improve the generalization capability of transformer backbones. As shown in Table V, we apply our method OSDFD to various models, including ViT [17] and Swin [85]. The results demonstrate that OSDFD consistently enhances performance across all backbones, leading to significant improvements in AUC scores in the cross-manipulation setting. Notably, the highest performance gains are observed in ViT-L (+0.167 AVG AUC) and Swin-L (+0.112 AVG AUC), suggesting that OSDFD is particularly effective when integrated with larger backbones. While the average cross-manipulation performance improves with increasing model capacity, larger models tend to require significantly more computing resources and inference time, thereby impeding their real-world deployment. These findings indicate that OSDFD is a flexible and effective module for improving the generalizability of transformer-based face forgery detection models.
In this work, we adopt a Parameter-Efficient Fine-Tuning (PEFT) strategy, a reliable approach, to train the face forgery detection model. Specifically, we only optimize the Adapter and LoRA layers while keeping the other ViT parameters fixed with ImageNet weights. We compare the PEFT strategy with the full fine-tuning (FT) strategy to validate its effectiveness. Table VI presents the cross-manipulation (c23) AUC results, where DF, FF, FS, and NT denote the testing manipulation types. The models are trained on three manipulations and tested on the remaining one. The two experimental settings employ two training strategies while maintaining the same network architectures. The results in Table VI demonstrate that the PEFT approach outperforms the fully FT strategy across all experimental trials. Notably, the average AUC score of PEFT exhibits a substantial enhancement, increasing from 68.2% to 84.3%. Hence, it can be concluded that the improvements in face forgery detection performance benefit from the PEFT strategy adopted. Furthermore, the PEFT strategy effectively preserves the abundant ImageNet prior knowledge while learning the forgery-aware knowledge through the designed Adapter and LoRA modules in an adaptive manner, reinforcing the reliability of our research.
TABLE V: Cross-manipulation AUC performance on FF++ (c23) dataset with different backbones.
Backbone OSDFD DF $F$ FS NT AVG
ViT-T - 0.853 0.658 0.488 0.584 0.646
ViT-T 0.919 0.860 0.563 0.709 0.763
ViT-S - 0.870 0.688 0.465 0.566 0.647
ViT-S 0.929 0.897 0.600 0.791 0.804
ViT-L - 0.850 0.708 0.572 0.593 0.681
ViT-L 0.955 0.941 0.710 0.785 0.848
Swin-S - 0.897 0.768 0.400 0.666 0.683
Swin-S 0.922 0.867 0.538 0.759 0.772
Swin-B - 0.877 0.715 0.463 0.668 0.681
Swin-B 0.943 0.888 0.583 0.734 0.787
Swin-L - 0.923 0.743 0.446 0.646 0.690
Swin-L 0.956 0.906 0.601 0.745 0.802
TABLE VI: Cross-Manipulation AUC performance on FF++ (c23) dataset with different training strategies.
Strategy DF $F$ FS NT AVG
Fully FT 0.896 0.731 0.496 0.603 0.682
PEFT (Ours) 0.962 0.932 0.698 0.781 0.843
Figure 9: Average cross-manipulation ACC scores across four experimental trials with respect to different mixed feature ratios.
Figure 10: Cross-dataset evaluation results using different seeds: (a) AUC; (b) EER.
Figure 11: Grad-CAM [86] visualization results of the ViT-B baseline and OSDFD on six unforeseen Deepfake datasets.
IV-E2 Effectiveness of FSM
To further evaluate the effectiveness of the FSM module, we report additional metrics, including mAP, ACC, and FNR, in Table VII. The cross-dataset evaluation results with and without FSM are presented in Table VII, showing that the inclusion of FSM consistently improves face forgery detection performance. Notably, FSM significantly enhances the FNR, indicating that it effectively mitigates forgery domain gaps between training and testing sets. This, in turn, enhances the model’s ability to detect forgery faces from another point of view.
IV-E3 Effectiveness of CDC
To validate the effectiveness of the CDC operator, we conduct an ablation study by integrating various conventional convolutions with different kernel sizes. Table VIII presents the model’s cross-dataset experimental results using different local feature extractors, where Conv3×3, Conv5×5, and Conv7×7 represent standard convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, respectively. The Second-order Difference Convolution (SDC) module, as proposed by Fei et al. [54], employs a second-order local anomaly learning mechanism to detect anomalies in local regions. These baseline results are obtained by replacing the CDC convolution in Fig. 4 with the corresponding operators. The results demonstrate that the CDC operator consistently outperforms other local feature extractors across various unforeseen datasets, which can be attributed to its outstanding capability for local high-frequency feature extraction.
IV-E4 Impacts of different ratios of mixed features
Our proposed method utilizes a feature mixture module to augment forgery domains in the feature space. As illustrated in Fig. 9, the $x$-axis represents the ratio of the mixed forgery features among all forgery features. We analyze the impact of this ratio by plotting the cross-manipulation (c23) ACC scores against ratio values in Fig. 9. Similarly, we present the average score across four cross-manipulation evaluation trials. A ratio of ‘0.0’ indicates the pure use of mixed forgery features during training, while ‘1.0’ means no forgery-style mixture is applied. Notably, relying solely on original or mixed forgery features results in limited face forgery detection accuracy. In contrast, combining original and mixed forgery features enhances the model’s generalization capability by augmenting the source forgery domains. The model achieves the highest average ACC score when the ratio of mixed forgery features is 50%. This observation underscores the effectiveness of the designed forgery feature mixture module from another perspective.
IV-F Visualization results
In this subsection, we provide visualization results to illustrate further the effectiveness of the proposed face forgery detection method.
IV-F1 T-SNE feature map visualization
In Fig. 12, we present the t-SNE [87] feature embedding visualization results of both the ViT-B baseline and our method OSDFD. Fig. 12 (a) and (b) showcase the results on the unseen CDF dataset, while Fig. 12 (c)-(h) depict the results on the WDF, DFDC, and FFIW datasets, respectively. Red and blue dots represent real and fake data samples in these visualizations. The manifolds of real and fake feature distributions entangle for the baseline model. This observation indicates that the baseline model struggles to discriminate forgery faces from the real ones. In contrast, OSDFD effectively separates real and fake images in the feature space. Specifically, for CDF and FFIW datasets, it is obvious that OSDFD can better discriminate real faces from fake ones compared to the ViT baseline. The AUC score goes from 72.35 to 83.35 on the CDF dataset, and 68.49 to 84.61 on the FFIW dataset. In turn, ViT-B can achieve good detection performance on WDF and DFDC datasets, while our proposed method OSDFD, benefiting from the designed forgery-aware parameter efficient fine-tuning and the forgery style mixture modules, exhibits a better face forgery detection performance.
Figure 12: t-SNE feature embedding visualization of (a) the ViT-B baseline on the CDF dataset, (b) OSDFD on the CDF dataset, (c) the ViT-B baseline on the WDF dataset, (d) OSDFD on the WDF dataset, (e) the ViT-B baseline on the DFDC dataset, (f) OSDFD on the DFDC dataset, (g) the ViT-B baseline on the FFIW dataset, and (h) OSDFD on the FFIW dataset. The red and blue dots represent real and fake data samples, respectively.
TABLE VII: Impacts of Forgery Style Mixture (FSM).
Setting CDF DFR WDF
[…]
w/o FSM 88.91 72.87 30.69 93.42 84.28 15.40 84.74 70.94 23.61
w/ FSM 90.33 75.05 22.82 95.25 88.22 11.01 86.70 72.77 16.13
IV-F2 Grad-CAM map visualization
We visualize the Grad-CAM [86] maps to identify attention regions of the developed face forgery detector for input queries, as shown in Fig. 11. These maps illustrate the results of the ViT-B baseline model and OSDFD across six unforeseen Deepfake datasets: CDF, WDF, DFDC, DFDC-P, DFR, and FFIW. These datasets encompass practical environmental scenarios, including lighting conditions, head poses, image quality, and facial occlusion. Compared to the baseline model, our proposed method accurately highlights informative manipulated regions. The baseline model sometimes focuses on irrelevant non-forgery areas, and these ‘non-interesting’ regions can lead to wrong decisions. Conversely, our method effectively identifies regions with significant forgery artifacts, contributing to superior face forgery detection performance. Specifically, our approach consistently demonstrates outstanding gradient-based localization across diverse environmental variables, such as extreme head poses in WDF, abnormal color hues in DFDC, facial occlusion (glasses) in DFDC-P, and extreme facial area ratios in FFIW. These observations further underscore the outstanding generalizability of our method in real-world applications.
TABLE VIII: Comparison with other local-frequency extractor.
Setting CDF DFR WDF AVG
[…]
Conv3×3 82.26 26.66 94.57 11.59 78.00 29.20 84.94 22.48
Conv5×5 80.39 27.32 91.35 15.75 76.98 30.18 82.91 24.42
Conv7×7 80.44 27.71 92.92 14.25 78.10 29.09 83.82 23.68
SDC [54] 79.14 27.89 92.87 11.21 77.72 28.32 83.24 22.47
CDC 83.35 24.47 95.54 9.90 78.53 28.46 85.81 20.94
IV-F3 Cross-dataset results using different seeds
To mitigate the influence of different seeds, we train our models for seven trials using seeds 0-6. The cross-dataset AUC and EER performances are shown in Fig. 10 (a) and (b), respectively. The results indicate that our approach consistently achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations. This robustness further underscores the superiority of our method from an additional perspective.
V Conclusions and Future Work
This paper introduces a parameter-efficient and generalized approach towards open-set face forgery detection. Our method, OSDFD, incorporates lightweight LoRA and Adapter layers into ViT backbones, updating only these modules during training to learn global and local forgery-aware features efficiently. On the other hand, we propose a novel forgery-style mixture module that augments the forgery source domains. In this vein, our method achieves SOTA generalizability and robustness in face forgery detection. An ablation study further underscores that the designed network modules and objective functions contribute to the final detection performance. Our proposed method, OSDFD, can be flexibly adapted to other backbones in a plug-and-play manner. Further ablation experiments demonstrate that the adopted parameter-efficient training strategy is effective in face forgery detection. Our proposed method will shed light on developing future face forgery detectors for real-world applications.
While our proposed method is effective, we limit our scope herein to non-sequential Deepfake detection. Adapting our approach to leverage abundant sequential information could significantly improve video-level face forgery detection. The designed components can be seamlessly integrated into video transformers. Extending our designed method to video transformers for video-level Deepfake detection opens an important research path forward. We anticipate that the proposed modules can well adapt to video ViTs and enhance the generalizability. Furthermore, our method exhibits strong robustness against common image perturb ations, as shown in Fig. 8. However, its performance degrades under the highest perturbation severity (level 5). Additionally, although our approach contributes to safeguarding digital media integrity, Deepfake detection models may exhibit bias across different demographic groups due to imbalances in the training dataset. In future work, our aim is to improve robustness and ensure fairness in Deepfake detection.